OpenStudy (anonymous):

If you simplify x • 5x by multiplying like terms, your answer is 5 times x raised to some exponent. What is the exponent of x? (Points : 3)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Not one.

OpenStudy (skullpatrol):

$$\Huge 25=5 \cdot5=5^2$$

OpenStudy (skullpatrol):

The number 25 can be written as 5*5 and is called a power of 5.

OpenStudy (skullpatrol):

Second power of 5: $$\Huge 5^2=5 \cdot5$$

OpenStudy (skullpatrol):

$$\Huge x^2= ?$$

OpenStudy (anonymous):

\[x \times x\]

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

what about this... \[x^m \times x^n=x^{m+n}\]... also any number without exponent is always a number raise to 1 such as \[a^1=a\]